# Boosting Robustness by Clipping Gradients in Distributed Learning

### v1

@@ 6 Empirical Evaluation
SOTA robust aggregations fail in extreme heterogeneity. Figure 3b shows that for all values of $f\in\{1,...,5\}$, Robust-DSGD with CWTM $\circ$ NNM and GM $\circ$ NNM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%. This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity. However, composing them with ARC increases the breakdown point of these aggregations to $f=3$. Indeed, for $f=1$ and 2, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to achieve accuracies greater than 85% in the worst case. However, when $f\geq 3$, their performance degrades, although CWTM $\circ$ NNM $\circ$ ARC reaches a satisfactory accuracy close to 70%. Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system. In Figure 1 of Section 1 where $\alpha=0.5$, the performances of ARC and no clipping are comparable for $f\leq 3$ for both aggregations.
Figure 4: FOE [43] on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
However, when $f\geq 5$, the improvement induced by ARC is much more visible. Particularly when $f=7$, ARC enables CWTM $\circ$ NNM and GM $\circ$ NNM to reach accuracies greater than 80% in the worst case, whereas the same aggregations without clipping yield accuracies below 40%. This observation confirms the raise of the breakdown point caused by ARC to robust aggregations in practice. Plots for $\alpha=0.1$ and 1 convey the same observations (see Appendix E.1).
Improved robustness on CIFAR-10. We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$. Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations. For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a large variance across seeds). For instance, GM $\circ$ NNM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC. In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance). On the contrary, executing the same methods without clipping deteriorates the performance of Robust-DSGD, with GM $\circ$ NNM and CWMed $\circ$ NNM reaching 16% and 13.7% in worst-case accuracy, respectively. This can also be seen in Figure 4, where FOE completely degrades the learning when ARC is not used, despite CWTM $\circ$ NNM reaching a peak of 40% at the start of the learning.
7 Conclusion
We have introduced a pre-aggregation clipping scheme, named adaptive robust clipping (ARC), that helps to circumvent the general limitations on the robustness to adversarial workers in distributed learning, under the standard heterogeneity model of $(G,B)$-gradient dissimilarity. We have obtained formal guarantees on the improvement in the learning error of state-of-the-art robust distributed gradient descent (Robust-DGD) when incorporating ARC. We have validated our theoretical findings through an exhaustive set of experiments on benchmark image classification tasks.
@@ 7 Conclusion
The algorithm starts by computing the norms of all input vectors, which takes $\mathcal{O}(nd)$ time. Note that this step is also required for static clipping. The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time. Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select [22] which has an average-case complexity of $\mathcal{O}(n)$. Finally, clipping the vectors requires $\mathcal{O}(nd)$ time. Therefore, the total complexity of ARC is $\mathcal{O}(nd+n\log(n))$. Overall, the complexity is comparable to static clipping, given that the sorting (or search) is crucially dimension independent.
@@ Appendix B Proofs for Results in Section 4
Consider first the case $\kappa=0$. Since $\kappa\geq\frac{f}{n-2f}$  [^6] Proposition 6, [3], we have that $f=0$. Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$. Since $\kappa=0$, the output of $\mathbf{F}$ will correspond to the average of its input, which implies that
@@
Appendix C Proofs for Results in Section 5
Proofs for Theorem 5.1, Corollary 5.2, Theorem 5.4 and Corollary 5.5 are presented in appendices C.2C.3C.4 and C.5, respectively.
C.2 Proof of Theorem 5.1
Our proof relies on the following lemma.

### v3

@@ 4 Empirical Evaluation
5 Improved Guarantee of Robust-DGD with ARC
In Section 4, we empirically demonstrated that ARC outperforms SOTA aggregation methods without clipping, despite these methods being theoretically proven to be optimal. This naturally raises the question: why does ARC, which shares similar convergence guarantees, significantly outperform plain Robust-DSGD? To address this, we now focus on the influence of model initialization on the robustness of aggregation methods, considering both theoretical results and empirical insights.
5.1 Improvement of convergence guarantees
In this section, we characterize the improvement of Robust-DGD with ARC over the lower bound recalled in Lemma 2.3. Specifically, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$. First, we establish in Lemma 5.1 a key property of ARC, which is crucial for deriving the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
Lemma 5.1 (Bounded Aggregation Output).
Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method. For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, the following holds true:
$\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (1)
This result indicates that incorporating ARC bounds the norm of the aggregated output by the largest norm of the honest gradients, i.e., $\max_{i\in\mathcal{H}}\left\lVert{x_{i}}\right\rVert$. In Appendix C, we provide the proof of Lemma 5.1 and further demonstrate in Lemma C.1 that this property is specific to ARC, i.e., classic $(f,\kappa)$-robust aggregation methods do not inherently exhibit this behavior. This property enables us to obtain the following improvement on the asymptotic convergence guarantee of Robust-DGD with ARC.
In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in 2.3. For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, let
$\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right),$
where $\rho$ denotes a real value. Recall from Lemma 2.3 that $\mathsf{BP}$ and $\varepsilon_{o}$ are the breakdown point and the lower bound on the stationarity error when $f/n<\mathsf{BP}$, respectively, under $(G,B)$-gradient dissimilarity. We obtain the following theorem for Robust-DGD with ARC. Let $\Delta_{o}$ be a real value such that $\mathcal{L}_{\mathcal{H}}({{\theta}_{1}})-\mathcal{L}_{\mathcal{H}}^{*}\leq\Delta_{o}$. The proof of Theorem 5.2 is deferred to Appendix C.
Theorem 5.2.
Suppose $B>0$ and there exists $\zeta\in\mathbb{R}^{+}$ such that $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$. Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$ $\circ\ \mathbf{NNM}$ , $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$ and $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$. Consider an arbitrary real value $\xi_{o}\in(0,1)$. Let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$. For any $\boldsymbol{\upsilon}\in(0,1)$, if $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $0<\xi\leq\min\left\{\frac{\boldsymbol{\upsilon}}{\Psi(G,B,\rho)},\xi_{o}\right\}$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\boldsymbol{\upsilon}~{}\varepsilon_{o}$.
Theorem 5.2 demonstrates that Robust-DGD with ARC improves over the lower bound $\varepsilon_{o}$ when the ratio $f/n$ is sufficiently close to the breakdown point ($\mathsf{BP}$), provided the local gradients at model initialization are bounded. This, in turn, leads to an improvement over the original learning guarantee of Robust-DGD (as stated in Lemma 2.6), which applies for arbitrary model initialization.
Influence of model initialization and data heterogeneity. Note that the smaller the bound on the initial gradients (i.e., the better the model initialization), the greater the improvement induced by ARC. Specifically, a decrease in $\zeta$ leads to a reduction in $\rho$, which subsequently lowers $\Psi(G,B,\rho)$. As a result, for a fixed value of $\xi$ (i.e., for a fixed ratio of adversarial workers), the condition $\xi\leq\boldsymbol{\upsilon}/\Psi(G,B,\rho)$ is satisfied for a smaller $\boldsymbol{\upsilon}$, thereby yielding a lower stationarity error. This theoretical deduction is empirically validated in Section 5.2. A similar improvement occurs when increasing data heterogeneity, while keeping all other factors unchanged. Specifically, as $G$ increases, $\Psi(G,B,\rho)$ decreases, allowing for a smaller reduction factor $\boldsymbol{\upsilon}$ and thus a larger improvement in performance. This trend is also validated empirically in Figures 3 and 4a.
Influence of $\boldsymbol{f/n}$. Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 2a and 4b. Indeed, we show in Theorem C.4 (a more complete version of Theorem 5.2) that ARC effectively increases the $\mathsf{BP}$ of ($f,\kappa$)-robust methods from $\frac{1}{2+B^{2}}$ to $\frac{1}{2}$, provided the initial honest gradients are bounded in norm. Lastly, we would like to recall that when the workers’ gradients at model initialization are arbitrarily large, the convergence guarantee of Robust-DGD with ARC reduces to that of classic Robust-DGD (see Section 3). In other words, there is no downside to incorporating ARC.
5.2 Influence of model initialization on empirical robustness
Figure 6: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, both with and without ARC, on MNIST with $\alpha=0.1$, 10 honest worker, and $f=1$ Byzantine worker. The x-axis represents settings with progressively worse model initialization. This figure complements Figure 2b.
Figures 2b and 6 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers. We investigate two regimes of heterogeneity: extreme heterogeneity and $\alpha=0.1$, and consider $f=1$ adversarial worker. The experiment proceeds as follows: we begin with the default model parameters initialized by PyTorch (i.e., the same ones used in Section 4), which represent a well-chosen set of initial parameters. These parameters are then scaled multiplicatively by a factor $\mu$ where larger values of $\mu$ correspond to progressively worse initialization, and vary $\mu\in\{1,...,5\}$. The results show that, under well-initialized conditions ($\mu=1$), ARC significantly enhances the performance of Robust-DSGD, achieving a substantial improvement in worst-case maximal accuracy, particularly under extreme heterogeneity. In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 2b). This ($\mu=1$) corresponds to the initialization conditions of the empirical results presented in Section 4. As $\mu$ increases (i.e., the initialization worsens), the performance of ARC-enhanced Robust-DSGD gradually declines, with noticeable degradation starting from $\mu=4$. Nevertheless, even at this point, ARC still offers a performance advantage over Robust-DSGD without clipping. By $\mu=5$, the performance of Robust-DSGD with ARC converges to that of plain Robust-DSGD, both achieving an accuracy of around 10%. Another key observation from Figures 2b and 6 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD. Both exhibit similarly poor performance when $\mu=5$, and their accuracies are comparable when $\mu\leq 3$, especially when $\alpha=0.1$. This suggests that ARC is particularly effective at exploiting good model initialization, similar to the performance of DSGD in the absence of Byzantine workers. In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 2b. These findings highlight the important influence of model initialization on the robustness of aggregation methods, and empirically validate our theoretical findings in Section 5.1.
6 Related Work
Gradient clipping is a well-known technique for tackling exploding gradients in deep learning (Goodfellow et al., 2016). It has been extensively analyzed in the centralized setting (Zhang et al., 2020b; a; Koloskova et al., 2023), with applications to differential privacy (Pichapati et al., 2019). Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise Gorbunov et al. (2020; 2022); Danilova (2023). The considered proof techniques, however, do not apply directly to our adversarial distributed learning setting. A similar study on the impact of clipping also exists for distributed settings (Zhang et al., 2022; Khirirat et al., 2023).
In the context of robust distributed learning, prior work has proposed iterative clipping for robust aggregation (Karimireddy et al., 2021). The clipping scheme, however, has not been shown to induce any improvement over other SOTA robustness schemes. Recent work (Malinovsky et al., 2023) has proposed pre-aggregation clipping using temporal gradient differences, in conjunction with variance reduction, to tackle partial worker participation when adversarial workers can form a majority. We, however, propose and study pre-aggregation clipping as a tool to improve the robustness obtained by a general class of aggregation rules. Other prior work (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a) that used static pre-aggregation clipping to enhance the empirical robustness of Robust-DGD, did not provide any formal explanation.
While adaptive clipping schemes, similar to ARC, have been proposed and studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted. The robustness guarantees in (Gupta & Vaidya, 2020; Liu et al., 2021) only apply to strongly convex loss functions, under a specific redundancy condition (namely, $2f$-redundancy). We consider non-convex loss functions, as well as a generic heterogeneous setting of $(G,B)$-gradient dissimilarity. In contrast to the clipping scheme proposed in Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f/n$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients. Moreover, we prove a deterministic robustness property of ARC (see Theorem 3.2), whereas Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024) only provide probability guarantees assuming the non-adversarial inputs to be i.i.d. with the distribution satisfying certain special properties.
7 Conclusion & Discussion
We introduced Adaptive Robust Clipping (ARC), a pre-aggregation clipping scheme designed to harness the empirical benefits of gradient clipping, while preserving the worst-case optimal convergence guarantees of Robust-DGD. Unlike existing adaptive clipping schemes, ARC does not require additional tuning since it only relies on the standard robustness parameter, i.e., the tolerable fraction of adversarial workers. Through theoretical analysis, we explained ARC’s ability to enhance the robustness of Robust-DGD, particularly when the model is well-initialized. This phenomenon was validated through comprehensive experiments on standard image classification datasets. In short, our experiments showed that ARC consistently boosts the performance of Robust-DGD, especially in scenarios with high heterogeneity and large fraction of adversarial workers.
Future direction. Our work also reveals a gap between theory and practice in Byzantine machine learning (ML). While Robust-DGD and ARC-enhanced Robust-DGD share the same worst-case convergence guarantees, their empirical performances are drastically different. In fact, and unsurprisingly, Byzantine ML theory focuses on worst-case guarantees, which, although essential, may not fully capture the practical realities of many learning settings. In practice, we often operate under favorable conditions (e.g., well-initialized models), where worst-case guarantees have limited relevance. This gap opens the door for future work that prioritizes practically-driven research in Byzantine ML, under conditions that are realizable in real-world scenarios. It encourages the development of robust distributed learning methods, like ARC, that can take full advantage of favorable practical conditions, thereby yielding relevant theoretical guarantees and superior empirical performance.
@@ 7 Conclusion & Discussion
The algorithm starts by computing the norms of all input vectors, which takes $\mathcal{O}(nd)$ time. Note that this step is also required for static clipping. The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time. Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select Hoare (1961) which has an average-case complexity of $\mathcal{O}(n)$. Finally, clipping the vectors requires $\mathcal{O}(nd)$ time. Therefore, the total complexity of ARC is $\mathcal{O}(nd+n\log(n))$. Overall, the complexity is comparable to static clipping, given that the sorting (or search) is crucially dimension independent.
@@ Appendix B Proofs for Results in Section 3
Consider first the case $\kappa=0$. Since $\kappa\geq\frac{f}{n-2f}$  [^5] Proposition 6, Allouah et al. (2023a), we have that $f=0$. Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$. Since $\kappa=0$, the output of $\mathbf{F}$ will correspond to the average of its input, which implies that
@@
Appendix C Proofs of the Results in Section 5
In this section, we first prove the Bounded Aggregation Output property of ARC in Lemma 5.1, and show that the other SOTA aggregation rules do not satisfy it in Lemma C.1. Then, to prove Theorem 5.2, we first prove Lemma C.2. We assume throughout this appendix that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$. This assumption can be made without loss of generality, see Appendix C.1.
See 5.1
Proof.
Without loss of generality and for the sake of simplicity, let us suppose that the vectors $x_{1},\dots,x_{n}$ are indexed such that $\left\lVert{x_{1}}\right\rVert\geq\left\lVert{x_{2}}\right\rVert\geq\dots\geq\left\lVert{x_{n}}\right\rVert$. Let $(\tilde{x}_{1},\dots,\tilde{x}_{n})=ARC{}(x_{1},\dots,x_{n})$ be the clipped version of these vectors after applying ARC.
Case 1: When $f=0$, $S=[n]$ and we clip all the vectors by $\left\lVert{x_{1}}\right\rVert=\max_{i\in S}\left\lVert{x_{i}}\right\rVert$, hence for any $j$, $\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$. Using the fact that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$, we have
$\left\lVert{\mathbf{F}\circ ARC{}(x_{1},\dots,x_{n})}\right\rVert=\left\lVert{\mathbf{F}(\tilde{x}_{1},\dots,\tilde{x}_{n})}\right\rVert\leq\max_{j\in[n]}\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$ (17)
Case 2: When $0<f<\frac{n}{2}$, as presented in Algorithm 2, ARC clips all the vectors $x_{1},\dots,x_{n}$ to $\left\lVert{x_{k+1}}\right\rVert$ where $k=\lfloor 2\frac{f}{n}(n-f)\rfloor$. Hence, for any $j\in[n]$,
$\left\lVert{\tilde{x}_{j}}\right\rVert\leq\left\lVert{x_{k+1}}\right\rVert\enspace.$ (18)
We have that $k+1=\lfloor 2\frac{f}{n}(n-f)\rfloor+1\geq 2\frac{f}{n}(n-f)$.
One can show that
$\frac{2f}{n}(n-f)-f$ $=f(1-\frac{2f}{n})$
$>0,\ \ \ \forall f\in\left]0,\frac{n}{2}\right[$ (19)
Hence, $\frac{2f}{n}(n-f)>f$, which gives $k+1>f$, and since $k+1$ is an integer, we have $k+1\geq f+1$. Hence we know that at least $f+1$ vectors have their norm greater or equal to the clipping threshold $C=\left\lVert{x_{k+1}}\right\rVert$,
$\underbrace{\left\lVert{x_{1}}\right\rVert\leq\dots\leq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\leq\underbrace{\left\lVert{x_{k+2}}\right\rVert\leq\dots\leq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
Hence, at most $n-f-1$ vectors will not be clipped. For any subset of $S\subset[n]$ of size $n-f$, at least one of the vectors has its norm equal to the clipping threshold (i.e. $\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{k+1}}\right\rVert$) or will be clipped (i.e. $\max_{i\in S}\left\lVert{x_{i}}\right\rVert>\left\lVert{x_{k+1}}\right\rVert$). Hence, we have that
$\max_{i\in S}\left\lVert{x_{i}}\right\rVert\geq\left\lVert{x_{k+1}}\right\rVert\enspace.$ (20)
Combining (18) and (20), we have that for all $j\in[n]$,
$\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$
Using the fact that for any set of vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(x_{1},\ldots,x_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{x_{i}}\right\rVert$, we have
$\left\lVert{\mathbf{F}\circ ARC{}(x_{1},\dots,x_{n})}\right\rVert=\left\lVert{\mathbf{F}(\tilde{x}_{1},\dots,\tilde{x}_{n})}\right\rVert\leq\max_{j\in[n]}\left\lVert{\tilde{x}_{j}}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert\enspace.$ (21)
Lemma C.1.
For any aggregation rule $\mathbf{F}\in\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK},\mathbf{CWTM}\circ\mathbf{NNM},\\ \mathbf{CWMed}\circ\mathbf{NNM},\mathbf{GM}\circ\mathbf{NNM},\mathbf{MK}\circ\mathbf{NNM}\}$, there exists $f\in[0,\frac{n}{2})$, a set of vectors $x_{1},\dots,x_{n}$ and a subset $S\subset[n]$, $|S|=n-f$ such that
$\mathbf{F}(x_{1},\dots,x_{n})\nleq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (22)
Proof.
Let us consider $(x_{1},x_{2},x_{3})=((0,1),(1,0),(1,1))$, and $S=\{x_{1},x_{2}\}$,
For $\mathbf{F}{}\in\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM}\}$, we have $\left\lVert{\mathbf{F}(x_{1},x_{2},x_{3})}\right\rVert=\left\lVert{(1,1)}\right\rVert=\sqrt{2}>\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{1}}\right\rVert=1$.
For $\mathbf{F}{}\in\{\mathbf{MK},\mathbf{CWTM}\circ\mathbf{NNM},\mathbf{CWMed}\circ\mathbf{NNM},\mathbf{GM}\circ\mathbf{NNM},\mathbf{MK}\circ\mathbf{NNM}\}$, we have $\left\lVert{\mathbf{F}(x_{1},x_{2},x_{3})}\right\rVert=\left\lVert{(0.5,1)}\right\rVert=\sqrt{1.25}>\max_{i\in S}\left\lVert{x_{i}}\right\rVert=\left\lVert{x_{1}}\right\rVert=1$. ∎
Lemma C.2.
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}L}{T}+5\kappa\left(G^{2}+B^{2}\rho^{2}\right).$
Our proof for Lemma C.2 relies on the following sub-result.